# FlexiCodec: A Dynamic Neural Audio Codec for Low Frame Rates

### v2

@@ Abstract
Neural audio codecs are foundational to speech language models. It is expected to have a low frame rate and decoupled semantic and acoustic information. A lower frame rate codec can reduce the computational cost of speech language models by shortening the sequence length. Recent studies have developed 12.5Hz low-frame-rate audio codecs, but even lower frame rate codecs remain underexplored. We find that a major challenge for very low frame rate tokens is missing semantic information. This paper introduces FlexiCodec to address this limitation. FlexiCodec improves semantic preservation with a dynamic frame rate approach and introduces a novel architecture featuring an ASR feature-assisted dual stream encoding and Transformer bottlenecks. With dynamic frame rates, it uses less frames at information-sparse regions through adaptively merging semantically similar frames. A dynamic frame rate also allows FlexiCodec to support inference-time controllable frame rates between 3Hz and 12.5Hz. Experiments on 6.25Hz, 8.3Hz and 12.5Hz average frame rates confirm that FlexiCodec excels over baseline systems in semantic information preservation and delivers a high audio reconstruction quality. We also validate the effectiveness of FlexiCodec in language model-based TTS. Demos are available at: https://flexicodec.github.io.
@@ 1 Introduction
The neural audio codec technique, originally designed for waveform compression (Zeghidour et al., 2021; Défossez et al., 2022), is now widely used in various tasks including Language Model (LM)-based text-to-speech (TTS) (Wang et al., 2023; Li et al., 2024; Wang et al., 2024b; Du et al., 2024a; Wang et al., 2025b; Guo et al., 2024) and multimodal LLMs (Défossez et al., 2024; Zeng et al., 2024; Ding et al., 2025). By compressing raw speech into compact discrete tokens, neural audio codecs enable the application of auto-regressive LLM paradigms to the speech domain.
@@
However, a fundamental challenge arises from the high temporal resolution of these codecs. State-of-the-art models like EnCodec (Défossez et al., 2022), DAC (Kumar et al., 2024), and SpeechTokenizer (Zhang et al., 2023a) operate at frame rates exceeding 50Hz, representing one second of audio with over 50 tokens. This high density, compared to the $\sim$4.5Hz rate of typical text representations (Wang et al., 2024a), creates two major problems for AR models: (1) a significant computational burden due to the quadratic complexity of attention, and (2) a severe frame rate mismatch between text and audio modalities that may degrade LLM performance (Wang et al., 2024a).
@@
While previous works have proposed 12.5Hz solutions, a significant gap remains compared to the $\sim$4.5Hz frame rate of text. Furthermore, research into neural audio codecs operating below 12.5Hz is limited. Our initial experiments revealed that pushing to even lower frame rates leads to significant information loss, particularly the omission of semantic/phonetic content. Driven by this observation, we have the following motivations for designing our low frame rate codec: (1) Dynamic Frame Rate: A fixed low rate inevitably discards transient phonetic details, whereas a dynamic rate may adapt to the phonetic complexity to encode more details. (2) Richer Semantics: SSL features, trained for reconstruction, can be redundant. Features from an ASR model, trained for text prediction, may offer a more concentrated source of semantic information. (3) Controllable Frame Rate: Existing codecs typically operate at one or more fixed frame rates. A continuously controllable rate would allow users to dynamically trade off performance and efficiency for downstream tasks.
@@
  • We propose FlexiCodec  [^1] Reproduced models will be available at https://github.com/amphionspace/flexicodec. To our best knowledge, it is the one of first neural audio codecs under 10Hz frame rate, and the first work to explore dynamic frame rate on low-frame-rate neural audio codecs. We develop a novel ASR feature-assisted dual stream codec architecture with Transformer bottlenecks.
@@ 2.1 Low-Frame-Rate Neural Audio Codecs
Neural audio codecs convert continuous speech into discrete tokens. SoundStream (Zeghidour et al., 2021), Encodec (Défossez et al., 2022), and DAC (Kumar et al., 2024) focused on audio compression, relying on residual vector quantization (RVQ) Zeghidour et al. (2021) and operating at high bitrates ($\geq$4kbps) and high frame rates ($\geq$50Hz). WavTokenizer (Ji et al., 2024), TS3-Codec (Wu et al., 2024), SemantiCodec (Liu et al., 2024), and StableCodec (Parker et al., 2024) used a single VQ or FSQ (Mentzer et al., 2023) codebook. They delivered good audio quality at low bitrates (around 1kbps), but operate at a high frame rates ($\geq$40Hz).
@@ 2.2 Dynamic-Rate Compression of Images and Audios
In the audio domain, one research trend has been the syllable-level semantic unit discoveries which are inherently dynamic-rate. SD-HuBERT Cho et al. (2024) finetuned HuBERT (Hsu et al., 2021) with sentence-level self-distillation, and showed that its features distinguish syllable boundaries. SylBoost (Baade et al., 2024) can discover discrete syllabic units using a min-cut algorithm on the feature self-similarity matrix, followed by a k-means clustering. It experimented on 8.33Hz and 6.25Hz units. However, compared to neural audio codec tokens, syllable units only encode coarse semantic information, and necessitate external speech synthesis models to produce audio.
@@ 3 FlexiCodec: A Dynamic Low-Frame-Rate Neural Audio Codec
Semantic (RVQ-1) Quantization The dynamic-rate ASR features are quantized using a Finite Scalar Quantizer (FSQ) (Mentzer et al., 2023) to produce discrete semantic tokens (we denote as RVQ-1 tokens). FSQ projects the input representations $e_{s}$ into a $D$-dimensional low-rank space, where the value of each dimension is quantized using rounding operation ROUND into $L$ levels. The quantized low-rank vector $\bar{e}$ is subsequently projected back to the original dimension $\bar{e}$:
@@
The semantic token index $q_{s}$ is obtained by: $q_{s}=\sum_{j=0}^{D-1}\bar{e_{s}}L^{j}$. Additionally, we wrap the FSQ block with small ConvNeXt (Liu et al., 2022) blocks to increase its representation ability, and apply an L2 loss $L_{\text{feat}}$ to align the semantic token embeddings with the unquantized semantic features.
Acoustic (RVQ-rest) Quantization To encode the detailed acoustic information that is not captured in the semantic tokens, following Li et al. (2025a), we compute a residual by subtracting the dynamic-rate ASR feature from the dynamic-rate waveform feature. This residual is then quantized using an (N1)(N-1)-layer Residual Vector Quantization (RVQ) (Zeghidour et al., 2021) [^2] We have not applied FSQ for acoustic quantization because FSQ is a single-layer quantization, and we have not discovered a multi-layer FSQ practice in literature. module to produce acoustic tokens $q_{2:N}\in Z^{(N-1)\times\hat{T}}$, where $\hat{T}$ denotes the sequence length after frame merging. We employ quantizer dropout (Zeghidour et al., 2021) during training. That is, only the RVQ-1 to RVQ-$n$ layers are subsequently decoded, where $n\in[1,N]$ is randomly chosen. When $n=1$, only the semantic encoding stream is used.
@@ 4.2 Examining the Impact of Very Low Frame Rates
We first investigate the performance of representative audio codecs at very low frame rates. Since open-source baselines operating below 12.5Hz are unavailable, we created three new baseline versions by retraining DAC (Kumar et al., 2024) and DualCodec (Li et al., 2025a) to operate at 12.5Hz, 8.3Hz, and 6.25Hz, respectively. To adapt these systems for lower frame rates, we increased their encoder downsampling rates and enlarged their codebook sizes to be consistent with FlexiCodec. Detailed configurations are provided in Appendix E.1. Our findings are as follows.
@@ 4.3 Analysis and Ablation of Dynamic Frame Rate
$\bullet$ Dynamic frame rate improves semantic information preservation. Tables 4 and 4 compare FlexiCodec and its variants that are retrained with fixed frame rate (FFR). To obtain the FFR variants, we modify their codec encoder strides to output static 8.3Hz or 6.25Hz, and maintain the total parameter count (detailed in Appendix E.1). Table 4 shows that the FFR variants have consistently worse WERs than FlexiCodec. The gap is larger at the lower frame rate. Specifically, the 6.25Hz FFR variant increases RVQ-1 WER by a relative 26%, the ASR probing WER by 21%, and the RVQ1:8 WER by 8%. These results highlight that dynamic frame rate improves semantic information preservation by a mechanism of phonetic complexity-adaptive frame rate allocation. We also note that dynamic frame rate may perform even better on real-world data, which tend to have longer silence regions that are highly compressible. We leave this investigation as future work.
@@ 4.4 Comparison with Open-Source Codecs at Various Bitrates
In this experiment, we compare FlexiCodec with open-source neural audio codecs spanning various bitrate and frame rates. Information about the baseline codecs are provided in Appendix E. We categorize them into 3 bitrate classes, and FlexiCodec at 12.5Hz, 8.3Hz, and 6.25Hz average frame rates fall into each category, enabling bitrate-consistent comparisons. Table 5 presents the results.
@@
<<0.7kbps Acoustic Bitrate
@@ 6 Conclusion
In this work, we introduced FlexiCodec, a novel neural audio codec designed for very low frame rate operation. By incorporating a dynamic frame rate, an ASR-assisted dual-stream architecture, and Transformer-based frame merging/unmerging modules, FlexiCodec effectively preserves semantic information at low rates. Experiments confirm its strong performance in low frame rate and low bitrate speech coding, controllable frame rate, and effectiveness in a downstream TTS system. Limitaion and future work are discussed in Appendix H.
7 Reproducibility Statement
To ensure the reproducibility of our work, we provide comprehensive details on methodology, training configurations, and evaluation setup throughout the paper and appendices. All datasets used, including Librilight-Large, LibriSpeech, and TIMIT, are publicly available. The source code for FlexiCodec, along with scripts for data processing, training, and evaluation, is provided as supplementary material. We will release our code under an open-source license after the review period.
@@ B.2 TTS Experiments
We use the E2TTS (Eskimez et al., 2024) test suite [^4] https://github.com/microsoft/e2tts-test-suite to test our models. Its test data consists of 1,132 (reference, test) pairs from Librispeech-PC-test-clean (Meister et al., 2023). We evaluate both objective and subjective metrics. The objective metrics include word error rate (WER) and speaker similarity (SIM-o). We report the real-time factor (RTF) of each system tested on an RTX5000 GPU. RTF is calculated as the ratio of the total processing time to the total duration of generated audios. The WER metric uses a Hubert-Large-LS960-ft ASR system (Hsu et al., 2021), and the SIM-o calculates the cosine similarity of speech samples of each test utterance against reference clips using a WavLM-large-based speaker verification model. For subjective evaluations, we evaluate Naturalness Mean Opinion Score (NMOS) and Quality Mean Opinion Score (QMOS). For each MOS, six professional linguistic experts evaluate 10 audios per system. The human evaluation instructions are attached in Appendix J. Baseline systems include E2TTS (Eskimez et al., 2024), VALL-E (Wang et al., 2023), CosyVoice (Du et al., 2024a), and SparkTTS (Wang et al., 2025b), FireRedTTS (Guo et al., 2024).
@@
Interestingly, for the AR stage, lower frame rates do not always degrade performance. With 50Hz NAR, the 8.3Hz and 6.25Hz AR configurations achieve comparable or even slightly better WER (2.5 vs. 2.5 vs. 3.2), SIM-o (0.64 vs. 0.65 vs. 0.65), and MOS scores compared to 12.5Hz AR. To explain this phenomenon, we suggest that (1) lower frame rate tokens might have better semantic and acoustic information disentanglement, and (2) a shorter sequence may alleviate the challenge of modeling long-range dependencies, allowing the Transformer to learn more effective attention patterns.

### v3

@@ Abstract
Neural audio codecs are foundational to speech language models. It is expected to have a low frame rate and decoupled semantic and acoustic information. A lower frame rate codec can reduce the computational cost of speech language models by shortening the sequence length. Recent studies have developed 12.5Hz low-frame-rate audio codecs, but even lower frame rate codecs remain underexplored. We find that pushing existing audio codecs to very low frame rates loses much semantic information. We suggest that low-frame-rate codecs’ limitations are in both insufficient semantic decoupling and insufficient time resolution at capturing transient phonetic details. This paper introduces FlexiCodec to address these limitations. FlexiCodec improves semantic preservation with a dynamic frame rate approach and introduces a novel architecture featuring an ASR feature-assisted dual stream encoding and Transformer bottlenecks. With dynamic frame rates, it uses less frames at information-sparse regions through adaptively merging semantically similar frames. A dynamic frame rate also allows FlexiCodec to support inference-time controllable frame rates between 3Hz and 12.5Hz. Experiments on 6.25Hz, 8.3Hz and 12.5Hz average frame rates confirm that FlexiCodec excels over baseline systems in semantic information preservation and delivers a high audio reconstruction quality. We also validate the effectiveness of FlexiCodec in language model-based TTS. Audio demos are available at: https://flexicodec.github.io. Code is available at: https://github.com/amphionteam/flexicodec.
@@ 1 Introduction
The neural audio codec technique, originally designed for waveform compression (Zeghidour et al., 2021; Défossez et al., 2022), is now widely used in various tasks including Language Model (LM)-based text-to-speech (TTS) (Wang et al., 2023; Li et al., 2024; Anastassiou et al., 2024; Du et al., 2024a; Wang et al., 2025b; Guo et al., 2024) and multimodal LLMs (Défossez et al., 2024; Zeng et al., 2024a; Ding et al., 2025). By compressing raw speech into compact discrete tokens, neural audio codecs enable the application of auto-regressive LLM paradigms to the speech domain.
@@
However, a fundamental challenge arises from the high temporal resolution of these codecs. State-of-the-art models like EnCodec (Défossez et al., 2022), DAC (Kumar et al., 2024), and SpeechTokenizer (Zhang et al., 2023a) operate at frame rates exceeding 50Hz, representing one second of audio with over 50 tokens. This high density, compared to the $\sim$4.5Hz rate of typical text representations (Wang et al., 2024), creates two major problems for AR models: (1) a significant computational burden due to the quadratic complexity of attention, and (2) a severe frame rate mismatch between text and audio modalities that may degrade LLM performance (Zeng et al., 2024b; Wang et al., 2024).
@@
While previous works have proposed 12.5Hz solutions, a significant gap remains compared to the $\sim$4.5Hz frame rate of text. Furthermore, research into neural audio codecs operating below 12.5Hz is limited. Our initial experiments revealed that pushing existing codecs to below 12.5Hz leads to significant loss of semantic/phonetic content. We suggest two reasons for this result. First, at very low frame rates, the limited information capacity forces a trade-off between acoustic fidelity and semantic representation. Existing codecs may be primarily limited by an insufficient decoupling of semantics from acoustics. This is exemplified by recent work on syllable-level unit discovery (Cho et al., 2024a; Baade et al., 2024), which successfully extracts semantic units at 5-8Hz but largely discards the fine-grained acoustic details, prosody, and timing required for high-fidelity reconstruction. We discuss these works in Section 2.2. Second, we propose that current fixed-rate downsampling can lead to the loss of transient phonetic details, whereas natural speech units like syllables and phonemes are inherently dynamic in their rate of occurrence. Driven by these observations, we have the following motivations for designing our low frame rate codec: (1) Dynamic Frame Rate: A fixed low rate inevitably discards transient phonetic details, whereas a dynamic rate may adapt to the phonetic complexity to encode more details. (2) Richer Semantics: SSL features, trained for reconstruction, can be redundant. Features from an ASR model, trained for text prediction, may offer a more concentrated source of semantic information. (3) Controllable Frame Rate: Existing codecs typically operate at one or more fixed frame rates. A continuously controllable rate would allow users to dynamically trade off performance and efficiency for downstream tasks.
@@
  • We propose FlexiCodec, a novel codec that pushes the boundaries of low-frame-rate audio tokenization. To our knowledge, it is among the first codecs to achieve high-quality, reconstructible audio at average frame rates below 10Hz. We also introduce a novel method for dynamic frame rate allocation in the low-frame-rate regime (3-12.5Hz), guided by ASR features. A review of recent below 10Hz codecs and dynamic-rate codecs are provided in Sections 2.1 and 2.2.
@@ 2.1 Low-Frame-Rate Neural Audio Codecs
Neural audio codecs convert continuous speech into discrete tokens. SoundStream (Zeghidour et al., 2021), Encodec (Défossez et al., 2022), and DAC (Kumar et al., 2024) focused on audio compression, relying on residual vector quantization (RVQ) (Zeghidour et al., 2021) and operating at high bitrates ($\geq$4kbps) and high frame rates ($\geq$50Hz). WavTokenizer (Ji et al., 2024), TS3-Codec (Wu et al., 2024), SemantiCodec (Liu et al., 2024), and StableCodec (Parker et al., 2024) used a single VQ or FSQ (Mentzer et al., 2023) codebook. They delivered good audio quality at low bitrates (around 1kbps), but operate at a high frame rates ($\geq$40Hz).
@@ 2.2 Dynamic-Rate Compression of Images and Audios
In the audio domain, one research trend has been the syllable-level semantic unit discoveries which are inherently dynamic-rate. SD-HuBERT (Cho et al., 2024b) finetuned HuBERT (Hsu et al., 2021) with sentence-level self-distillation, and showed that its features distinguish syllable boundaries. Sylber (Cho et al., 2024a) leveraged SD-HuBERT as pseudo labels, and trained a student model with syllable encoding ability at around 5 syllables per second. SylBoost (Baade et al., 2024) can discover discrete syllabic units using a min-cut algorithm on the feature self-similarity matrix, followed by a k-means clustering. It experimented on 8.33Hz and 6.25Hz units. However, compared to neural audio codec tokens, syllable units only encode coarse semantic information, and necessitate external speech synthesis models to produce audio.
@@ 3 FlexiCodec: A Dynamic Low-Frame-Rate Neural Audio Codec
Semantic (RVQ-1) Quantization The dynamic-rate ASR features are quantized using a Finite Scalar Quantizer (FSQ) (Mentzer et al., 2023) to produce discrete semantic tokens (we denote as RVQ-1 tokens). FSQ projects the input representations $e_{s}$ into a $D$-dimensional low-rank space, where the value of each dimension is quantized using rounding operation ROUND into $L$ levels. The quantized low-rank vector $\bar{e}$ is subsequently projected back to the original dimension $\hat{e}$:
@@
Straight-through estimation (Bengio et al., 2013) is employed to estimate the gradients in the ROUND function. The semantic token index $q_{s}$ is obtained by: $q_{s}=\sum_{j=0}^{D-1}\bar{e_{s}}L^{j}$. Additionally, we wrap the FSQ block with small ConvNeXt (Liu et al., 2022) blocks to increase its representation ability, and apply an L2 loss $L_{\text{feat}}$ to align the semantic token embeddings with the unquantized semantic features.
Acoustic (RVQ-rest) Quantization To encode the detailed acoustic information that is not captured in the semantic tokens, following Li et al. (2025a), we compute a residual by subtracting the dynamic-rate ASR feature from the dynamic-rate waveform feature. This residual is then quantized using an (N1)(N-1)-layer Residual Vector Quantization (RVQ) (Zeghidour et al., 2021) [^1] We have not applied FSQ for acoustic quantization because FSQ is a single-layer quantization. We consider the integration of more advanced multi-layer FSQ (rFSQ) schemes, e.g., (Parker et al., 2024), to be a promising direction for future improvements. module to produce acoustic tokens $q_{2:N}\in Z^{(N-1)\times\hat{T}}$, where $\hat{T}$ denotes the sequence length after frame merging. We employ quantizer dropout (Zeghidour et al., 2021) during training. That is, only the RVQ-1 to RVQ-$n$ layers are subsequently decoded, where $n\in[1,N]$ is randomly chosen. When $n=1$, only the semantic encoding stream is used. Straight-through estimation (Bengio et al., 2013) is used to backpropagate through the codebook lookup.
@@ 4.2 Examining the Impact of Very Low Frame Rates
We first investigate the performance of representative audio codecs at very low frame rates. Since open-source baselines operating below 12.5Hz are unavailable, we created three new baseline versions by retraining DAC (Kumar et al., 2024) and DualCodec (Li et al., 2025a) to operate at 12.5Hz, 8.3Hz, and 6.25Hz, respectively. To adapt these systems for lower frame rates, we increased their encoder downsampling rates and enlarged their codebook sizes to be consistent with FlexiCodec. Detailed configurations are provided in Appendix G.1. Our findings are as follows.
@@ 4.3 Analysis and Ablation of Dynamic Frame Rate
$\bullet$ Dynamic frame rate improves semantic information preservation. Tables 4 and 4 compare FlexiCodec and its variants that are retrained with fixed frame rate (FFR). To obtain the FFR variants, we modify their codec encoder strides to output static 8.3Hz or 6.25Hz, and maintain the total parameter count (detailed in Appendix G.1). Table 4 shows that the FFR variants have consistently worse WERs than FlexiCodec. The gap is larger at the lower frame rate. Specifically, the 6.25Hz FFR variant increases RVQ-1 WER by a relative 26%, the ASR probing WER by 21%, and the RVQ1:8 WER by 8%. These results highlight that dynamic frame rate improves semantic information preservation by a mechanism of phonetic complexity-adaptive frame rate allocation. We also note that dynamic frame rate may perform even better on real-world data, which tend to have longer silence regions that are highly compressible. We leave this investigation as future work.
@@ 4.4 Comparison with Open-Source Codecs at Various Bitrates
In this experiment, we compare FlexiCodec with open-source neural audio codecs spanning various bitrate  [^3] Bitrate indicates the amount of data stored or transmitted per unit of time, measured in kilobits per second (kbps). It can be calculated by ($\log_{2}$ Codebook_size) $\times$ Frame_rate $\times$ Num_rvq_layer. A lower bitrate codec has higher compression ratio but usually lower audio quality. Since storing and transmitting each of FlexiCodec’s frame length attribute takes $\log_{2}8=3$ bits, we have added $3\timesframe$_rate to FlexiCodec’s bitrate in Table 5. and frame rates. Information about the baseline codecs are provided in Appendix G. We categorize them into 3 bitrate classes, and FlexiCodec at 12.5Hz, 8.3Hz, and 6.25Hz average frame rates fall into each category, enabling bitrate-consistent comparisons. Table 5 presents the results.
@@
$\leq$0.7kbps Acoustic Bitrate
TAAE-25Hz-700bps - 0.70 / 2q 953M - 4.19 2.84 4.33 4.98 0.58
TAAE-25Hz-400bps 0.40 0.40 / 1q 953M 5.21 5.21 2.61 4.34 4.10 0.53
SemantiCodec-50Hz 0.68 0.68 / 1q 921M 4.97 4.97 2.38 3.03 4.94 0.63
SemantiCodec-25Hz 0.34 0.34 / 1q 1878M 23.8 23.8 1.89 2.93 5.92 0.40
@@ 5 More Experiments
$\bullet$ Codec efficiency analysis. In Appendix E, we benchmark the encoding and decoding speed of FlexiCodec. The results show that FlexiCodec has a low Real-Time Factor (RTF) of 0.018 for encoding and 0.006 for decoding across all frame rates. This shows that our model is a practical and efficient solution for integration into larger pipelines like TTS systems.
$\bullet$ Generalization to out-of-domain and multilingual data. In Appendix F, we evaluate the robustness of our English-trained model on the diverse, out-of-domain Emilia (He et al., 2024) dataset. The results show that (1) FlexiCodec maintains strong semantic and acoustic performance on out-of-domain English speech, confirming it is not overfitted to audiobook data. (2) While the semantic tokens struggle on unseen languages in a zero-shot setting, the full acoustic reconstruction remains competitive with multilingual models. (3) A fine-tuning on an unseen language, Chinese, can adapt FlexiCodec’s semantic tokens and the merging scheme to the new language.
@@ 6 Conclusion
In this work, we introduced FlexiCodec, a novel neural audio codec designed for very low frame rate operation. By incorporating a dynamic frame rate, an ASR-assisted dual-stream architecture, and Transformer-based frame merging/unmerging modules, FlexiCodec effectively preserves semantic information at low rates. Experiments confirm its strong performance in low frame rate and low bitrate speech coding, controllable frame rate, and effectiveness in a downstream TTS system. Limitaion and future work are discussed in Appendix J.
7 Acknowledgements
This work is supported by the NSFC Grant 62376237; the Shenzhen Science and Technology Program ZDSYS20230626091302006; 2023 Shenzhen Stability Science Program; Program for Guangdong Introducing Innovative and Entrepreneurial Teams 2023ZT10X044.
@@ B.2 TTS Experiments
We use the E2TTS (Eskimez et al., 2024) test suite [^4] https://github.com/microsoft/e2tts-test-suite to test our models. Its test data consists of 1,132 (reference, test) pairs from Librispeech-PC-test-clean (Meister et al., 2023). We evaluate both objective and subjective metrics. The objective metrics include word error rate (WER) and speaker similarity (SIM-o). We report the real-time factor (RTF) of each system tested on an RTX5000 GPU. RTF is calculated as the ratio of the total processing time to the total duration of generated audios. The WER metric uses a Hubert-Large-LS960-ft ASR system (Hsu et al., 2021), and the SIM-o calculates the cosine similarity of speech samples of each test utterance against reference clips using a WavLM-large-based speaker verification model. For subjective evaluations, we evaluate Naturalness Mean Opinion Score (NMOS) and Quality Mean Opinion Score (QMOS). For each MOS, six professional linguistic experts evaluate 10 audios per system. The human evaluation instructions are attached in Appendix L. Baseline systems include E2TTS (Eskimez et al., 2024), VALL-E (Wang et al., 2023), CosyVoice (Du et al., 2024a), and SparkTTS (Wang et al., 2025b), FireRedTTS (Guo et al., 2024).
@@
Interestingly, for the AR stage, lower frame rates do not always degrade performance. With 50Hz NAR, the 8.3Hz and 6.25Hz AR configurations achieve comparable or even slightly better WER (2.5 vs. 2.5 vs. 3.2), SIM-o (0.64 vs. 0.65 vs. 0.65) compared to 12.5Hz AR. To explain this phenomenon, we suggest that (1) lower frame rate tokens might have better semantic and acoustic information disentanglement, and (2) a shorter sequence may alleviate the challenge of modeling long-range dependencies, allowing the Transformer to learn more effective attention patterns.